If 10a+2b=15 and 5a=4b,what is the value of a+b?

Answer by Alan3354(69443)

If 10a+2b=15 and 5a=4b,what is the value of a+b?
----------------
10a+2b=15
5a=4b ---> 10a = 8b
====================
10a+2b=15
10a = 8b
----
8b + 2b = 15
b = 1.5
-----
a = 1.2
===============
a+b = 2.7